According to the IRS (Internal Revenue Service, www.irs.gov), a single taxpayer with a taxable income over $319,100 in 2004 paid $92,592.50 plus 35% of the amount over $319,100. If Glen paid $255,836 in federal income tax for 2004, then what was his taxable income for 2004?
Answer by gonzo(654)

proglem:
According to the IRS (Internal Revenue Service, www.irs.gov), a single taxpayer with a taxable income over $319,100 in 2004 paid $92,592.50 plus 35% of the amount over $319,100. If Glen paid $255,836 in federal income tax for 2004, then what was his taxable income for 2004?
-----
let T = tax glen paid.
let I = glen's taxable income
general formula is T = $92592.5 + (.35*(I - $319100)
this becomes
$255836 = $92592.5 + (.35*(I - $319100)
subtracting 92592.5 from both side of the equation and it becomes
$163243.5 = (.35*(I - $319100)
removing parentheses from right side of equation and it becomes
$163243.5 = .35*I - .35*$319100
multiplying .35*$319100 and adding it to both sides of the equation and it becomes
$163243.5 + $111685 = .35*I
doing the addition on the left side of the equation and dividing both sides of the equation by .35 and it becomes
$274928.5/.35 = I
solving for I and the equation becomes
$785510 = I
-----
answer looks like Glen's income is $785510
to prove correct, calculate Glen's taxes.
formula again is T = 92592.5 + (.35*(I - $319100)
this becomes
T = 92592.5 + (.35*($785510 - $319100)
this becomes
T = 92592.5 + (.35*$466410)
this becomes
T = 92592.5 + $163243.5
which becomes
T = $255836
-----
answer is confirmed.
Glen's Taxable Income is $785,510.00